OpenStudy (anonymous):

. Perform the indicated operation and, if possible, simplify. Please show all of your work. -21/50 Divide by 9/14

OpenStudy (mertsj):

\[\frac{-21}{50}\div\frac{9}{14}=\frac{-21}{50}\times\frac{14}{9}\]

OpenStudy (mertsj):

Cancel 3 from the -21 and 9 Cancel 2 from the 14 and 50

OpenStudy (mertsj):

\[\frac{-7}{25}\times\frac{7}{3}=\frac{-49}{75}\]

OpenStudy (anonymous):

Thank You